handheld fiber optic meters with white light polarization interferometry are useful for measuring temperature, pressure, and strain in electrically noisy environments. the fixed costs associated with manufacturing are $750000 per year. if variable costs are 290 per unit and the company sells 4200 units per year, at what selling price per unit will the company break even?
Answer #1

Breakeven formula = Fixed Cost/(Price-Variable Cost)

4200 = 750000/(Price-290)

(Price-290) = 750000/4200 = 178.57

Price = 178.57+290 = 468.57

So, the price per unit is $468.57